11. If the income elasticity of demand for a product is 1.0, and if consumers’ income increases by 10%, the proportion of income spent on this product ………….. [1] P = 6.67 and Q = 300. [2] P = 20 and Q = 300. [3] P = 3 and Q = 275. [4] P = 2 and Q = 500


1
Expert's answer
2021-05-24T13:55:58-0400

Answer:

Since YED = 1.0, 10% increase in income (Y) results in 10% increase in demand (Q) i.e. Unitary elasticity. The proportion of income spent on:

(1) P= 6.67 and Q = 300;

"\therefore" New demand (Q) = 300 + (0.1"\times"300) = 300 + 30 = 330

Income spent = P "\times" Q = 6.67 "\times" 330 = 2,201.10


(2) P = 20 and Q = 300

New demand (Q) = 300 + (0.1"\times"300) = 300 + 30 = 330

Income spent = P "\times" Q = 20 "\times" 330 = 6,600


(3) P = 3 and Q = 275:

New demand (Q) = 275 + (0.1"\times"275) = 275 + 27.5 = 302.5

Income spent = P "\times" Q = 3 "\times" 302.5 = 907.50


(4) P = 2 and Q = 500

New demand (Q) =500 + (0.1"\times"500) = 500 + 50 = 550

Income spent = P "\times" Q = 2 "\times" 550 = 1,100